OpenStudy (anonymous):

What is the value of 6*12+4[(20*-6)+8]? Plz help

OpenStudy (anonymous):

if you don't get what it is saying here........ http://static.k12.com/bank_packages/files/media/mathml_69f0bfadc0ef250ac091608c33a55b7cb1862b4f_1.gif

OpenStudy (anonymous):

thats the equation

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

First, subtract 20 - 6.

OpenStudy (anonymous):

do you mind if i write this down to study?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

20-6=14. correct?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Then add 14 + 8.

OpenStudy (anonymous):

14+8=22

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Then multiply 6 x 12.

OpenStudy (anonymous):

6 x 12=72

OpenStudy (anonymous):

Yes. Now multiply 4 x 22.

OpenStudy (anonymous):

4 x 22=88

OpenStudy (anonymous):

Yes, so now, all you have to do is add 72 + 88.

OpenStudy (anonymous):

72+88=160

OpenStudy (anonymous):

Yes! So 160 is your answer! :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem. :)

OpenStudy (anonymous):

i am going to exit this question but can i tag you?